OpenStudy (anonymous):

Three less than twice the third of three consecutive integers equals the sum of the three integers. Find the integers...... Where do i even start? 1+2+3=(3*2-3)?

OpenStudy (ranga):

Let the three consecutive integers be: N, N + 1, N + 2 "Three less than twice the third of three consecutive integers equals the sum of the three integers." 2(N+2) - 3 = N + N + 1 + N + 2 solve for N.

OpenStudy (anonymous):

i see, thank you very much

OpenStudy (ranga):

You are welcome. What are you getting for N?

OpenStudy (anonymous):

i got -2

OpenStudy (ranga):

Yes, the three integers are: -2, -1 and 0.

OpenStudy (anonymous):

hey, how would i simplify \[\sqrt{4-x^2}\]

OpenStudy (ranga):

a^2 - b^2 = (a+b)(a-b) 4 - x^2 = 2^2 - x^2 = (2+x)(2-x) sqrt(4-x^2) = sqrt((2+x)(2-x))

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

can i jus take the twos out from there?